OpenStudy (anonymous):

2B SQUARE+23B+15/3B+8= WHAT IS THE ANSWER

OpenStudy (anonymous):

do u mean \[2b^2 +23b +(15\div3b) + 8\] sorry it was just a little hard to see

OpenStudy (anonymous):

2b square + 23b +15 /3b+8 =

OpenStudy (anonymous):

please use ( ) , [ ] or { }

OpenStudy (anonymous):

2b square + 23b +15 divide 3b +8 =?

OpenStudy (anonymous):

Would u please help me this problem. tHANKS

OpenStudy (anonymous):

do you want to know the roots, right?

OpenStudy (anonymous):

so 2b^2 + 23b +15 /(3b+8) = 0